# A stochastic network approach to clustering and visualising single-cell genomic count data

### v4

@@ Methods
The stochastic blockmodel [holland1983stochastic, karrer2011stochastic, qin2013regularized] is another latent position model, under which community detection is carried out by spectral clustering [newman2006modularity, qin2013regularized] and modularity maximisation [newman2006modularity]. Communities in a network are groups of highly interconnected nodes in that network (Figure 2a); ordering the nodes by community membership along the margins of the adjacency matrix reveals the blockmodel structure and modular nature of the network (Figure 2b). Very importantly, Newman equated spectral clustering of network nodes with maximising an estimator called the network modularity [newman2006modularity], unifying these different perspectives on equivalent methodologies. Network modularity is of particular relevance to genomic data, because modularity maximisation with a symmetric adjacency matrix $\mathbf{A}$ is the basis of the popular, reliable, and effective ‘Louvain clustering’ algorithm [blondel2008fast]. This symmetric $\mathbf{A}\in\{0,1\}^{n\times n}$ can be generated e.g., via a $k$-nearest neighbours graph derived from the data-matrix, in contrast to our first-principles model of the data-matrix $\mathbf{A}=\mathbf{X}\in\mathbb{Z}^{p\times n}_{\geq 0}$. Other ways to generate a symmetric adjacency matrix from genomic data matrix $\mathbf{X}$ include thresholding the sample correlation matrix, based on prior knowledge or principled transformations [bartlett2017network].
@@
To summarise, in the results that we present in Section 3 we focus on GMM-LE clustering for one of the node-types described (cells), i.e., clustering the rows of $\mathbf{V}\in\mathbb{R}^{n\times d}$. We visualise these nodes/cells in $\mathbf{Y}^{\top}\in\mathbb{R}^{n\times 2}$ with UMAP-LE, where each sample-point $j\in\{1,...,n\}$ in the visualisation can be coloured as in Figures 4, 5, and 6. Nodes of only one type (i.e., cells) are plotted as dots in those figures, but each of these nodes/dots is then coloured based on its network connections to the other type of node (i.e., genes). Specifically, the colour intensity of the dot plotted for each node (cell) $j$ is proportional to the number of connections $\sum_{i\in g}X_{ij}$ in the network $\mathbf{A}=\mathbf{X}\in\mathbb{Z}^{p\times n}_{\geq 0}$ that node $j$ has to a predefined set of markers $i\in g$ of the other node-type (genes). We also use a machine-learning model to identify a cancer-predictive component $l\in\{1,...d\}$ of the $d$-dimensional LE via its associative effect in the latent positions $\phi^{(R)}_{j}$, $j\in\{1,...,n\}$, estimated by $V_{l,j}$. Then we validate a cancer biomarker based on $U_{l,i}$, where $i\in g^{\prime}$ is from a set of genes $g^{\prime}$ chosen by similar reasoning to the discussion of cases 1 and 2 above.
@@
Under the RDPG and GRDPG models, we expect asymptotically multivariate-Gaussian distributions of data-points when transformed to their latent positions [athreya2016limit, tang2018limit, rubin2022statistical]. Hence, we cluster cells using a Gaussian mixture model (GMM) in the Laplacian eigenspace (LE) $\mathbf{V}\in\mathbb{R}^{n\times K^{\prime}}$ which we refer to as GMM-LE clustering, as follows. We note that a popular choice of $K^{\prime}$ is $K^{\prime}=K-1$ components, where $K$ is the number of clusters sought [riolo2014first].
@@
where $f_{\mathcal{N}}$ denotes Gaussian density. We note that clusters inferred using a Gaussian mixture model with diagonal covariance matrices (i.e., diagonal $\boldsymbol{\Sigma}_{k}$, $k\in\{1,...,K\}$) are equivalent to the clusters obtained from the K-means algorithm.
@@ Results
Figure 4, Figure 5, and Figure 6 show clustering and visualisation based on the Laplacian eigenspace (LE) latent positions estimated for the data-points/cells of, respectively, the human cortical development data-set, the human embryonic development data-set, and the breast cancer at-risk data-set. UMAP is used to project the n data-points into two dimensions for visualisation, based on their transformed latent positions in the estimated Laplacian eigenspace (UMAP-LE). GMM-LE clustering is also carried out by fitting a Gaussian mixture model in the same estimated LE. In the plots in these figures, each dot or data-point represents one cell (sample). In the cluster plots, the colours indicate the GMM-LE cluster that the cell is assigned to. In the remaining plots, the colours indicate the average log-expression levels of sets of marker genes defined by domain experts for specific cell-types that are relevant to each data-set. These sets of marker genes have been used previously by us in other studies relating to human neural development [bartlett2021two], human embryonic development [alanis2023mica], and breast cancer at-risk [bartlett2021inference].
@@
Research on human embryos is highly restricted, due to ethical and regulatory considerations. This constrains sample sizes in generated data-sets, restricting inferences that can be obtained from these samples. If a novel clustering method is able to increase the number of cells obtained from an embryo that can be reliably identified as a specific cell-type of interest in the embryo such as epiblast cells, then it would be potentially very valuable to biological science researchers because it would increase the effective sample size for these precious cells that can be used for further research based on those data. Epiblast cells are precursors to the ectoderm layer, one of three primary germ layers of the embryo; epiblast cells go on to form neural and epithelial cells [wainwright2010langman]. Using GMM-LE clustering, we were able to increase the number of cells that can be reliably identified as epiblast from a well-known study of human embryos [petropoulos2016single], compared to the latest research in the field [stirparo2018integrated]. To set the number of clusters K shown in Figure 5, we increased K until the epiblast cells formed a distinct cluster as shown by the UMAP projection of the Laplacian eigenspace (UMAP-LE), setting $K=8$. In Figure 5, cluster 5 identifies 68 epiblast cells according to expression of the marker genes NANOG and KLF17, and the absence of expression of GATA3 and SOX17 [blakeley2015defining]. These 68 cells include 44 of 45 cells that were previously identified as epiblast, as well as very importantly 24 cells newly identified as epiblast cells here. The full list of all 68 epiblast cells are given in Table S1.
@@
The main cell-types of the epithelial lineage in breast tissue are luminal progenitor cells and mature luminal cells (thought to be the cells-of-origin of, respectively, hormone-receptor negative and positive breast cancers) as well as basal cells [tharmapalan2019mammary]. As luminal progenitor cells are thought to be the cell-of-origin of highly aggressive triple negative breast cancers (TNBC), it is important to be able to identify these cells as accurately as possible, e.g., for quantifying their numbers in tissue samples from at-risk individuals. Important applications of accurately quantifying luminal progenitor cells include developing novel biomarkers and surrogate end-points for clinical trials of chemopreventative medicines [bartlett2022antiprogestins]. We have previously shown that GMM-LE clustering improves identification of luminal progenitor cells in single-cell genomic data compared to Seurat-Louvain clustering [bartlett2021inference]. In this study we identify the breast epithelial subtypes basal, luminal progenitor, and mature luminal cells in a different recently published data-set [pal2021single], again showing an improvement over Seurat-Louvain clustering (Table 2). The breast epithelial subtypes can be separated well in the UMAP projection of the data-points in the Laplacian eigenspace (UMAP-LE), and they can be separated well by GMM-LE clustering, as illustrated by Figure 6. To carry out GMM-LE clustering here, we again increased the number of clusters K until the breast epithelial subtypes were well separated visually, choosing $K=8$. We also note that plotting the cells with positions according to their principal component scores for the top two dimensions (i.e., the first two rows of $\mathbf{V}$, as in Figure S2) is less effective for visualising these cell types separately than generating a two-dimensional projection from the top 25 dimensions of $\mathbf{V}$.
@@
Predicting long-term patient survival outcome
As a further application of the proposed methodology, we developed a novel way to discover biomarkers of risk of aggressive cancer. We did this by testing the association of each dimension of the LE with a known genetic marker of predisposition to the most aggressive breast cancers, a mutation in the BRCA1 (FANCS) gene [tharmapalan2019mammary]. Individuals with a mutation in the BRCA1 (FANCS) gene have a much increased risk of breast cancer during their lifetime [kuchenbaecker2017evaluation]. This means that cells from the healthy tissue of BRCA1 mutation carriers may have some characteristics that predispose some of these cells to ultimately transform into aggressive cancer [bartlett2022antiprogestins, williams2024luminal]. These characteristics of cancer predisposition would also be expected to be reflected in the gene expression profiles of these cells. Therefore, we compared scores derived from each dimension of the LE latent positions in healthy cells from BRCA1 mutation carriers (BRCAmut) with those from cells from BRCA1 wild-type (WT). To do this, we fitted a generalised linear model with logit link-function, to the top 25 RSVs $\mathbf{V}\in\mathbb{R}^{n\times 25}$ as covariates (predictors), with BRCA1 mutation status of the donor of the cell sample as the response. The absolute value of the gradient of the fitted linear model for each of these dimension is shown in Figure 7a (with z-test p-values); dimension 12 had the largest absolute gradient and smallest p-value, so was chosen for the LE biomarker.
@@ Conclusions
In this paper we have presented a novel first-principles method of modelling single-cell genomic count data, by modelling these data as a bipartite multi-edge network. This view of single-cell genomic data opens up a wealth of new modelling approaches for this type of data based on stochastic networks theory. It also gives theoretical validation to well-established and highly successful methods for clustering these data based on community detection, such as Louvain clustering [blondel2008fast]. In particular, this view shows the promise of studying the eigenspace of the graph-Laplacian (the ‘Laplacian eigenspace’, LE), to identify and represent known and novel cell-types. Based on this, we have proposed a new method of visualising single-cell genomic data based on the UMAP projection of the representation of the data-points in the Laplacian eigenspace (UMAP-LE), and we have proposed GMM-LE clustering, or ‘Gaussian mixture modelling in the Laplacian eigenspace’, for genomic data. We show that UMAP-LE visualisations correspond to existing biological knowledge in several data-sets. We have also identified a component of the LE that is associated with predisposition towards aggressive breast cancer, and have validated a biomarker extracted from this LE component, showing significant association with long-term patient survival outcome.
@@
There are several outstanding questions still to address in this research. In particular, choosing the number of clusters for a particular data-set is a challenging problem, which we have addressed heuristically for each data-set here. Developing an automatic method for choosing the number of clusters for GMM-LE clustering is a priority. A straightforward approach to this problem is to set the number of clusters as one greater than the dimensionality of the Laplacian eigenspace that is used [riolo2014first]. We note that the latent dimensionality of the eigenspace of the graph-Laplacian can be estimated manually by identifying an eigengap in the ordered plot of singular values, or by using existing theory based on scree plots [zhu2006automatic]. Recent findings from stochastic network modelling provides refined methodology to choose the optimum number of clusters to estimate in the Laplacian eigenspace via the Bayesian Information Criterion (BIC) [priebe2019two]. Building on existing theory for bootstrapping network data with latent structure [levin2019bootstrapping] will enable estimates of uncertainty to be calculated for estimators of the number of latent clusters, as well as quantifying uncertainty on the estimated Laplacian eigenspace. Finally, we have previously proposed a novel method of estimating a latent space in which different cell-types or clusters can be separated well, based on the Mahalanobis distance [bartlett2021inference]. It will be interesting to investigate how this relates to the estimate of the Laplacian eigenspace, and refines it for identifying different cell-types from single-cell genomic data.
In summary, our proposed methodological advances in this work provide mathematical technologies that have the potential for large impact in biomedical science. This is achieved by providing theoretical justification for and improvement on existing and widely adopted methodology for identifying important cell-types in single-cell genomic data.
@@ Acknowledgements
The work carried out by TE Bartlett was partly supported by MRC grant (MR/P014070/1). The authors acknowledge the use of the UCL High Performance Computing Facilities, and associated support services, in the completion of this work. The results published herein are in whole or in part based upon data generated by the TCGA Research Network: https://www.cancer.gov/tcga
@@
Figure S2: The top two Laplacian eigenspace dimensions as two-dimensional latent positions, for the breast-cancer at-risk data-set. The top two rows of $\mathbf{V}$ display cells $j\in\{1,...,n\}$ coloured according to mean expression levels of validating marker genes for cell-types of interest..
Figure S3: Laplacian eigenspace / latent position dimensions as PC (principal component) scores, for the breast-cancer at-risk data-set. UMAP projections from the Laplacian eigenspace (UMAP-LE) display cells $j\in\{1,...,n\}$ coloured according to their RSV value $V_{j,l}$, for each dimension $l\in\{1,...,d\}$. Green and red indicate positive and negative values of $V_{j,l}$ (respectively), with colour intensity proportional to magnitude.

### v5

@@ Methods
The stochastic blockmodel [holland1983stochastic, karrer2011stochastic, qin2013regularized] is another latent position model, under which community detection is carried out by spectral clustering [newman2006modularity, qin2013regularized] and modularity maximisation [newman2006modularity]. Communities in a network are groups of highly interconnected nodes in that network (Figure 2a); ordering the nodes by community membership along the margins of the adjacency matrix reveals the blockmodel structure and modular nature of the network (Figure 2b). Very importantly, Newman equated spectral clustering of network nodes with maximising an estimator called the network modularity [newman2006modularity], unifying these different perspectives on equivalent methodologies. Network modularity is of particular relevance to genomic data, because modularity maximisation with a symmetric adjacency matrix $\mathbf{A}$ is the basis of the popular, reliable, and effective ‘Louvain clustering’ algorithm [blondel2008fast], as implemented in the Seurat package which is widely used in genomics [satija2015spatial]. This symmetric $\mathbf{A}\in\{0,1\}^{n\times n}$ is generated in Seurat-Louvain clustering from a k-nearest-neighbours graph derived from the data-matrix, in contrast to our first-principles model of the data-matrix $\mathbf{A}=\mathbf{X}\in\mathbb{Z}^{p\times n}_{\geq 0}$. Other ways to generate a symmetric adjacency matrix from genomic data matrix $\mathbf{X}$ include thresholding the sample correlation matrix, based on prior knowledge or principled transformations [bartlett2017network].
@@
To summarise, in the results that we present in Section 3 we focus on GMM-LE clustering for one of the node-types described (cells), i.e., clustering the rows of $\mathbf{V}\in\mathbb{R}^{n\times d}$. We visualise these nodes (cells) in $\mathbf{Y}^{\top}\in\mathbb{R}^{n\times 2}$ with UMAP-LE, where each sample-point $j\in\{1,...,n\}$ in the visualisation can be coloured as in Figures 4, 5, and 6. Nodes of only one type (i.e., cells) are plotted as dots in those figures. Each of these nodes/dots is then coloured based on the number of network edges that connect it to the other type of node (i.e., genes). Hence, the colour intensity of the dot plotted for each node (cell) $j$ is proportional to the number of connections $\sum_{i\in g}X_{ij}$ in the network $\mathbf{A}=\mathbf{X}\in\mathbb{Z}^{p\times n}_{\geq 0}$ that node $j$ has to a predefined set of markers $i\in g$ of the other node-type (genes). We also use a machine-learning model to identify a cancer-predictive component $l\in\{1,...d\}$ of the $d$-dimensional LE via its assortative effect in the transformations of latent positions $\phi^{(R)}_{j}$, $j\in\{1,...,n\}$, as estimated by $V_{l,j}$. Then we validate a cancer biomarker based on $U_{l,i}$, where $i\in g^{\prime}$ is from a set of genes $g^{\prime}$ chosen by similar reasoning to the discussion of cases 1 and 2 above.
@@
Under the RDPG and GRDPG models, we expect asymptotically multivariate-Gaussian distributions of data-points in their transformed latent positions [athreya2016limit, tang2018limit, rubin2022statistical]. Hence, we cluster cells using a Gaussian mixture model (GMM) in the Laplacian eigenspace (LE) $\mathbf{V}\in\mathbb{R}^{n\times K^{\prime}}$ which we refer to as GMM-LE clustering, as follows. We note that a popular choice of $K^{\prime}$ is $K^{\prime}=K-1$ components, where $K$ is the number of clusters sought [riolo2014first].
@@
where $f_{\mathcal{N}}$ denotes Gaussian density. We note that clusters inferred using a Gaussian mixture model with diagonal covariance matrices (i.e., diagonal $\boldsymbol{\Sigma}_{k}$, $k\in\{1,...,K\}$) would be equivalent to the clusters obtained from the K-means algorithm.
@@ Results
Figure 4, Figure 5, and Figure 6 show clustering and visualisation based on the Laplacian eigenspace (LE) transformed latent positions estimated for the data-points/cells of, respectively, the human cortical development data-set, the human embryonic development data-set, and the breast cancer at-risk data-set. UMAP is used to project the n data-points into two dimensions for visualisation, based on their transformed latent positions in the estimated Laplacian eigenspace (UMAP-LE). GMM-LE clustering is also carried out by fitting a Gaussian mixture model in the same estimated LE. In the plots in these figures, each dot or data-point represents one cell (sample). In the cluster plots, the colours indicate the GMM-LE cluster that the cell is assigned to. In the remaining plots, the colours indicate the average log-expression levels of sets of marker genes defined by domain experts for specific cell-types that are relevant to each data-set. These sets of marker genes have been used previously by us in other studies relating to human neural development [bartlett2021two], human embryonic development [alanis2023mica], and breast cancer at-risk [bartlett2021inference].
@@
Research on human embryos is highly restricted, due to ethical and regulatory considerations. This constrains sample sizes in generated data-sets, restricting inferences that can be obtained from these samples. If a novel clustering method is able to increase the number of cells obtained from an embryo that can be reliably identified as a specific cell-type of interest in the embryo such as epiblast cells, then it would be potentially very valuable to biological science researchers because it would increase the effective sample size for these precious cells that can be used for further research based on those data. Epiblast cells are precursors to the ectoderm layer, one of three primary germ layers of the embryo; epiblast cells go on to form neural and epithelial cells [wainwright2010langman]. Using GMM-LE clustering, we are able to increase the number of cells that can be reliably identified as epiblasts from a well-known study of human embryos [petropoulos2016single], compared to the latest research in the field [stirparo2018integrated]. To set the number of clusters K shown in Figure 5, we increased K until the epiblast cells formed a distinct cluster as shown by the UMAP projection of the Laplacian eigenspace (UMAP-LE), setting $K=8$. In Figure 5, cluster 5 identifies 68 epiblast cells according to expression of the marker genes NANOG and KLF17, and the absence of expression of GATA3 and SOX17 [blakeley2015defining]. These 68 cells include 44 of 45 cells that were previously identified as epiblast, as well as very importantly 24 cells newly identified as epiblast cells by our proposed methodology. The full list of all 68 epiblast cells are given in Table S2.
@@
The main cell-types of the epithelial lineage in breast tissue are luminal progenitor cells and mature luminal cells (thought to be the cells-of-origin of, respectively, hormone-receptor negative and positive breast cancers) as well as basal cells [tharmapalan2019mammary]. As luminal progenitor cells are thought to be the cell-of-origin of highly aggressive triple negative breast cancers (TNBC), it is important to be able to identify these cells as accurately as possible, e.g., for quantifying their numbers in tissue samples from at-risk individuals. Important applications of accurately quantifying luminal progenitor cells include developing novel biomarkers and surrogate end-points for clinical trials of chemopreventative medicines [bartlett2022antiprogestins].
@@
We have previously shown that GMM-LE clustering improves identification of luminal progenitor cells in single-cell genomic data compared to Seurat-Louvain clustering [bartlett2021inference]. In this study we identify the breast epithelial subtypes basal, luminal progenitor, and mature luminal cells in a different recently published data-set [pal2021single], again showing an improvement over Seurat-Louvain clustering (Table 2). As an additional comparison with Louvain clustering, we cluster these cells using the Louvain implementation available in the scikit-net package for Python [bonald2020scikit], again taking the genomic data-matrix as a bipartite network adjacency matrix. However, this approach finds 17 clusters, of which one is a giant cluster (comprising 97% of the nodes/cells), indicating that the k-nearest-neighbour graph construction step in the Seurat-Louvain implementation is important for using Louvain clustering with genomic count data. This result remained unchanged with either the default ‘Dugue’ or ‘Newman’ settings in the scikit-network-Louvain implementation.
@@
The breast epithelial subtypes can be separated well in the UMAP projection of the data-points in the Laplacian eigenspace (UMAP-LE), and they can be separated well by GMM-LE clustering, shown in Figure 6. To carry out GMM-LE clustering here, we again increase the number of clusters K until the breast epithelial subtypes are well separated visually, choosing $K=8$. We also note that plotting the cells with positions according to their principal component scores for the top two dimensions (i.e., the first two rows of $\mathbf{V}$, as in Figure S2) is less effective for visualising these cell types separately than generating a two-dimensional projection from the top 25 dimensions of $\mathbf{V}$.
Breast cancer at-risk dataset: predicting long-term patient survival outcome
As a further application of our proposed methodology, we develop a novel way to discover biomarkers of risk of aggressive cancer. We do so by testing the association of each dimension of the LE with a known genetic marker of predisposition to the most aggressive breast cancers, a mutation in the BRCA1 (FANCS) gene [tharmapalan2019mammary]. Individuals with a mutation in the BRCA1 (FANCS) gene have a much increased risk of breast cancer during their lifetime [kuchenbaecker2017evaluation]. This means that cells from the healthy tissue of BRCA1 mutation carriers may have some characteristics that predispose some of these cells to ultimately transform into aggressive cancer [bartlett2022antiprogestins, williams2024luminal]. These characteristics of cancer predisposition would also be expected to be reflected in the gene expression profiles of these cells. Therefore, we compare scores derived from each dimension of the LE transformed latent positions in healthy cells from BRCA1 mutation carriers (BRCAmut) with those from cells from BRCA1 wild-type (WT). To do so, we fit a generalised linear model with logit link-function to the top 25 RSVs $\mathbf{V}\in\mathbb{R}^{n\times 25}$ as covariates (predictors), with BRCA1 mutation status of the donor of the cell sample as the response. The absolute value of the gradient of the fitted linear model for each of these dimension is shown in Figure 7a (with z-test p-values); dimension 12 has the largest absolute gradient and smallest p-value, and hence is selected for the LE biomarker.
@@ Conclusions
In this paper we have presented a novel first-principles method of modelling single-cell genomic count data, by proposing to model the observed numbers of mRNA molecules in specific cells with a bipartite multi-edge network. This view of single-cell genomic data opens up a wealth of new modelling approaches for this type of data based on stochastic networks theory. Our proposed methodology is also closely related to well-established and highly successful methods for clustering these data based on community detection, such as Louvain clustering [blondel2008fast]. The work we have presented here shows the promise of studying the eigenspace of the graph-Laplacian (the ‘Laplacian eigenspace’, LE), to identify and represent known and novel cell-types. Based on this, we have proposed a new method of visualising single-cell genomic data based on the UMAP projection of the representation of the data-points in the Laplacian eigenspace (UMAP-LE), and we have proposed GMM-LE clustering, or ‘Gaussian mixture modelling in the Laplacian eigenspace’, for genomic data. We show that UMAP-LE visualisations correspond to existing biological knowledge in several data-sets. We have also identified a component of the LE that is associated with predisposition towards aggressive breast cancer, and have validated a biomarker extracted from this LE component, showing significant association with long-term patient survival outcome.
@@
There are several outstanding questions still to address in this research. In particular, choosing the number of clusters for a particular data-set is a challenging problem, which we have addressed heuristically for each data-set here. Developing an automatic method for choosing the number of clusters for GMM-LE clustering is a priority. A straightforward approach to this problem is to set the number of clusters as one greater than the dimensionality of the Laplacian eigenspace that is used [riolo2014first]. We note that the latent dimensionality of the eigenspace of the graph-Laplacian can be estimated manually by identifying an eigengap in the ordered plot of singular values, or by using existing theory based on scree plots [zhu2006automatic]. Recent findings related to stochastic network theory provide refined methodology to choose the optimum number of clusters to estimate in the Laplacian eigenspace via the Bayesian Information Criterion (BIC) [priebe2019two]. Building on existing theory for bootstrapping network data with latent structure [levin2019bootstrapping] will enable estimates of uncertainty to be calculated for estimators of the number of latent clusters, as well as quantifying uncertainty on the estimated Laplacian eigenspace. Finally, we note that we have previously proposed a novel method for estimating a latent space in which different cell-types or clusters can be separated well, based on the Mahalanobis distance [bartlett2021inference]. It will be interesting to investigate how that idea relates to the estimate of the Laplacian eigenspace, and how it may be refined for identifying different cell-types from single-cell genomic data.
In summary, the methodological advances that we present here provide mathematical technologies that have the potential for large impact in biomedical science. This is achieved by providing theoretical justification for and improvement on existing and widely adopted methodology for identifying important cell-types in single-cell genomic data.
@@ Acknowledgements
The work carried out by TE Bartlett was partly supported by MRC grant (MR/P014070/1). The authors acknowledge the use of the UCL High Performance Computing Facilities, and associated support services, in the completion of this work. The results published here are in part based on data generated by the TCGA Research Network: https://www.cancer.gov/tcga
@@
With FS Raw data
HCD 0.9998 0.8869
HED 0.9992 0.7340
BCAR 1.0000 0.8835
Table S1: Correlation coefficients comparing library size with the top RSV of the SVD of the graph-Laplacian. The top RSV corresponds to the largest singular value. Spearman correlation was used; all reported values correspond to $p<2.2\times 10^{-16}$ (asymptotic t-test approximation). Abbreviations: Feature selection (FS); Human cortical development (HCD); Human embryonic development (HED); Breast-cancer at-risk (BCAR).
@@
Figure S2: The top two Laplacian eigenspace dimensions as two-dimensional transformed latent positions, for the breast-cancer at-risk data-set. The top two rows of $\mathbf{V}$ display cells $j\in\{1,...,n\}$ coloured according to mean expression levels of validating marker genes for cell-types of interest..
Figure S3: Laplacian eigenspace / transformed latent position dimensions as PC (principal component) scores, for the breast-cancer at-risk data-set. UMAP projections from the Laplacian eigenspace (UMAP-LE) display cells $j\in\{1,...,n\}$ coloured according to their RSV value $V_{j,l}$, for each dimension $l\in\{1,...,d\}$. Green and red indicate positive and negative values of $V_{j,l}$ (respectively), with colour intensity proportional to magnitude.